OpenStudy (usukidoll):

Quick question what is the antiderivative of 3sinxcosx?

OpenStudy (usukidoll):

wouldn't my u be sinx and du cosx?

mathslover (mathslover):

Anti derivative .. hmn wait

mathslover (mathslover):

Can we say ; antiderivative as "integration" also?

OpenStudy (mayankdevnani):

@mathslover yes we can

mathslover (mathslover):

@.Sam. Please let me try?

OpenStudy (usukidoll):

I just need to know what the antiderivate of 3sinxcosx is and yes I am integrating from crud hold up let me find it. Got hazelnut on my fingers

mathslover (mathslover):

is it : \(\frac{-3}{2}cos^2x\) ?

mathslover (mathslover):

OH + C there also

sam (.sam.):

\[\int\limits_{}^{}3sinxcosx~dx \\ \\ Let~u=sinx \\ \\du=cosx~dx \\ \\ Then, \\ \\ \int\limits_{}^{}3u~du \\ \\ Integrate\]

mathslover (mathslover):

@UsukiDoll you were at right process.

sam (.sam.):

@mathslover oh sry didn't see your comment there.

mathslover (mathslover):

:) No problem @.Sam. , it is so nice of you to even say "sry" no mistake of yours' . Though, can you tell whether I am correct or not (scroll up to see my answer)

OpenStudy (mayankdevnani):

\[\int\limits_{0}^{\frac{\pi}{2}}\] 3cos(x)sin(x)dx=\[\frac{3}{2}\]

OpenStudy (usukidoll):

I need the antiderivative of 3sinxcosx integrating from pi/2 to 0. umm that's what I did. like u sinx du cosx and then it's 3/2 u^2 but for some reason my manual has it at 3/2 sin2t. ummm WHAT!

OpenStudy (mayankdevnani):

@UsukiDoll ok got it

OpenStudy (mayankdevnani):

@mathslover am i right

OpenStudy (usukidoll):

sin^2t maybe but sin2t? what the heck

mathslover (mathslover):

==lol

sam (.sam.):

@mathslover yes you are

OpenStudy (anonymous):

why not just u=sinx so du=cosxdx ???

OpenStudy (usukidoll):

ummm that's what I did bytemr

OpenStudy (usukidoll):

oops yes a =0 b = pi/2

OpenStudy (anonymous):

lol.... you said my name...

OpenStudy (usukidoll):

loloolol

OpenStudy (usukidoll):

I could've sworn it was like this... 

hartnn (hartnn):

3sinx cos x = 3/2 sin 2x does this ring the bell ?

OpenStudy (usukidoll):

Why is it sinx 2x?!?!?!?!?!?!?!!?

OpenStudy (usukidoll):

is there an identity that I missed @hartnn

hartnn (hartnn):

double angle formula: sin 2x = 2 sin x cos x cos 2x = cos^2 x- sin^2 x = 2cos^2x -1 = 1-2sin^2 x

OpenStudy (usukidoll):

yeah but that 2sinxcosx not 3sinxcosx and why is the 2 below the 3 in this one?

hartnn (hartnn):

@mathslover you were typing something ? 3sin x cos x = 3/2 (2sin x cos x) = 3/2 sin 2x

hartnn (hartnn):

its antiderivative will be -3/4 cos 2x

OpenStudy (usukidoll):

again what I don't understand is why the 3/2? the problem is 3sinxcosx. There is no 2sinxcosx. so what Do i DO?

hartnn (hartnn):

so u did not understand this ? 3sin x cos x = 3/2 (2sin x cos x) i just multiplied and divided by 2

OpenStudy (usukidoll):

ohhhhh..... -___-

OpenStudy (usukidoll):

that's legal? multiplying 2 and dividing 2?

hartnn (hartnn):

absolutely! because 2 is constant.

OpenStudy (usukidoll):

O______O!

OpenStudy (usukidoll):

ummm what math class is this normally in? Is it in trig?

hartnn (hartnn):

those double angle formulas? in trignometry..

OpenStudy (usukidoll):

no about the 3/2 2sinxcosx

hartnn (hartnn):

its general, no particular topic, u can do that anytime, anywhere...

OpenStudy (usukidoll):

THE HECK?

OpenStudy (usukidoll):

I got one more...so I'm opening a new question.

hartnn (hartnn):

ok.